# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v3

@@ Abstract
$C^{*}$-algebra-valued kernels could pave the way for the next generation of kernel machines. To further our fundamental understanding of learning with $C^{*}$-algebraic kernels, we propose a new class of positive definite kernels based on the spectral truncation. We focus on kernels whose inputs and outputs are vectors or functions and generalize typical kernels by introducing the noncommutativity of the products appearing in the kernels. The noncommutativity induces interactions along the data function domain. We show that it is a governing factor leading to performance enhancement: we can balance the representation power and the model complexity. We also propose a deep learning perspective to increase the representation capacity of spectral truncation kernels. The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels, addressing two of the foremost issues regarding learning in vector-valued RKHSs, namely the choice of the kernel and the computational cost.
@@ 1 Introduction
Kernel methods have been one of the most fundamental tools in machine learning (Schölkopf & Smola, 2001; Gretton et al., 2007; Hofmann et al., 2008; Muandet et al., 2017). They have been applied, for example, to ridge regression, principal component analysis, and support vector machine. Kernel methods are characterized by reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels. Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel. Product kernels, which are constructed by the product of multiple kernels, have also been considered (Schölkopf & Smola, 2001; Thomas, 2008).
@@
Standard positive definite kernels are scalar-valued, and are well-suited to learn scalar-valued functions. Kernel methods for vector- and function-valued outputs have also been investigated (Álvarez et al., 2012; Kadri et al., 2016). The kernels, in these cases, are instead operator-valued, and the associated feature space is vector-valued RKHSs (vvRKHSs) (Kadri et al., 2012, 2016; Minh et al., 2016). There are at least two challenges for vvRKHS methods: the computational cost and choice of kernels. A typical kernel is the separable kernel, which is defined by the product of a scalar-valued kernel and a positive semi-definite operator (Álvarez et al., 2012). Although applying the separable kernel is computationally efficient, it identifies dependencies between input and output variables separately, and cannot reflect information of input variables properly to output variables. Several attempts have been made to construct kernels that go beyond separable kernels. Huusari & Kadri (2021) proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement. Hashimoto et al. (2023a) proposed to use the product of circulant matrices and general squared matrices to construct kernels. Using this kernel, one can generalize the convolution and capture the effect of interactions of different Fourier components. However, if we need an $m$-dimensional vector-valued outputs with nonseparable kernels, then we have to construct an $mN$ by $mN$ Gram matrix, where $N$ is the sample size, and the computational cost is $O(m^{3}N^{3})$ in general. Thus, with vvRKHSs, we cannot resolve the two issues at the same time.
@@
In this work, we address these two challenges by introducing a new class of kernels based on the framework of reproducing kernel Hilbert $C^{*}$-module (RKHM) (Hashimoto et al., 2021). RKHM is a generalization of RKHS by means of $C^{*}$-algebra. $C^{*}$-algebra is a generalization of the space of complex values and has structures of the norm, product, and involution (Murphy, 1990; Lance, 1995). It unifies operators and functions. In this framework, kernels are generalized to $C^{*}$-algebra-valued kernel functions and allow us to consider function-valued kernels, leading to function-valued Gram matrices. By evaluating the values of the function-valued Gram matrix at $m$ different points, we obtain $m$ scalar-valued Gram matrices. This allows us to obtain an $m$-dimensional vector-valued outputs with the computational cost of $O(mN^{3})$, which alleviates the dependency on $m$ from cubic to linear compared to the case of vvRKHSs with nonseparable kernels.
@@
To obtain kernels going beyond the separable kernels, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra (D’Andrea et al., 2014; van Suijlekom, 2021; Connes & van Suijlekom, 2021). Our class of kernels is large and encompasses previous commutative kernel functions such as polynomial, product, and separable kernels. The proposed kernels are parameterized by a natural number $n$ corresponding to the dimension of the truncated space. They can be applied to both vector and functional inputs. For vector inputs, we regard the elements of them as the values of functions. We approximate the input functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices, whose $(i,j)$-entry depends only on $i-j$. Thus, $n$ describes the resolution of the discretization, and we call it the truncation parameter. We show that the proposed kernels converge to the commutative kernels as $n$ goes to infinity.
In the setting of functional data (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021), the proposed truncation kernels shed light on the good effects of discretization (setting $n$ as a finite number) on the learning process. The truncation parameter $n$ plays an important role from at least two perspectives. First, an important structure in typical commutative kernels is the product, and $n$ describes the noncommutativity of the product. For two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$. Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product. This situation corresponds to the case of $n=\infty$. On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$. This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$. This corresponds to the case $n<\infty$. Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity. We show that if $n$ is small, then the representation power is low, and the model complexity is small. On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large. The expectation loss for learning a function is controlled by the representation power and the complexity. By introducing the parameter $n$ and setting $n$ to balance them, we obtain higher performance compared to the typical commutative kernels. In addition, to resolve the issue of the interactions becoming weaker as $n$ becomes larger (i.e., representation power becomes higher), we propose a deep approach to obtain models with higher representation powers while keeping the interactions.
@@
  • We propose spectral truncation kernels, a new class of function-valued positive definite kernels that generalize typical kernels. The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter, which induces the interactions along the data function domain by introducing the noncommutativity into the learning process. (see Definition 6 and Subsection 3.1)
  • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels. The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity. (see Section 4)
  • We propose a deep approach to gain the representation power while keeping the noncommutativity in the kernels. (see Section 6)
Our framework allows us to go beyond the usual commutative kernels and to develop a new perspective on the significance of noncommutative structures in learning systems.
@@ 2.2 Spectral truncation on the torus
The product in $C(\mathbb{T})$ is commutative. However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a Toeplitz matrix, we can obtain a noncommutative product structure. Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$. Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square-integrable functions on $\mathbb{T}$. Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$. We approximate $M_{x}$ by $P_{n}M_{x}P_{n}$, i.e., by truncating the spectrum. Then, the $(j,l)$-entry of the representation matrix $R_{n}(x)\in\mathbb{C}^{n\times n}$ of $P_{n}M_{x}P_{n}$ is written as
@@
where $F_{n}(t)=(1/n)\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution. By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ (Brandolini & Travaglini, 1997; van Suijlekom, 2021).
@@
A generalization of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated (Travaglini, 1994; Brandolini & Travaglini, 1997). Let $P$ be a polyhedron and $jP=\{jz\,\mid\,z\in P\}$ for $j\in\mathbb{Z}$. For $t\in\mathbb{T}^{q}$, the Fejér kernel on $\mathbb{T}^{q}$ is defined as
@@ 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. To obtain vector- or functional-valued outputs, applying vvRKHSs has been investigated. However, as we stated in Section 1, to go beyond the separable kernel and reach higher performance, the computational cost scales as $O(m^{3}N^{3})$ for $m$-dimensional outputs with $N$ samples. To resolve this issue, we consider function-valued ($\mathcal{A}$-valued) kernels and RKHMs. As we will discuss in the last part of Section 5, the computational cost depends on $mN^{3}$ if we use an $\mathcal{A}$-valued kernel. Hashimoto et al. (2021) proposed to use the $\mathcal{A}$-valued kernel $k$ where $k(\cdot,\cdot)(z)$ is a complex-valued positive definite kernel for all $z\in\mathbb{T}$. Combining it with typical existing kernels (Hashimoto et al., 2023a; Schölkopf & Smola, 2001; Kadri et al., 2012; Álvarez et al., 2012), we obtain the following three kinds of $\mathcal{A}$-valued kernels.
@@
    product kernel $k^{\operatorname{prod},q}(x,y)(z)=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z))$,
  3. 3.
    separable kernel $k^{\operatorname{sep},q}(x,y)=\tilde{k}(x,y)\prod_{j=1}^{q}(a_{j}^{*})^{q}a_{j}^{q}$.
Here, $q\in\mathbb{N}$ is the degree of the products in the kernel, $\alpha_{i}\geq 0$ is the parameter of $k^{\operatorname{poly},q}$, and for $i=1,2$ and $j=1,\ldots,q$, $\tilde{k}_{i,j}:\mathbb{C}\times\mathbb{C}\to\mathbb{C}$ and $\tilde{k}:\mathcal{A}^{d}\times\mathcal{A}^{d}\to\mathbb{C}$ are complex-valued continuous positive definite kernels. In addition, $a_{j}\in\mathcal{A}$ is the parameter of $k^{\operatorname{sep},q}$.
We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the typical kernels in Example 2. The kernels in Example 2 involve the product of functions in $\mathcal{A}$. For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1). Thus, the kernels defined in Examples 2.1 and 2.2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$. If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship. On the other hand, if we introduce the truncation parameter $n$ and transform $x$ and $y$ into Toeplitz matrices using the map $R_{n}$ defined in Subsection 2.2, then the product of $R_{n}(x)$ and $R_{n}(y)$ is noncommutative. Focusing on this feature, we define $\mathcal{A}$-valued kernels based on Example 2, but they have interactions along $z\in\mathbb{T}$.
@@
$k_{n}^{\operatorname{prod},q}(x,y)$ $=S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(\tilde{k}_{1,j}(x,y))^{*}\prod_{j=1}^{q}R_{n}(\tilde{k}_{2,j}(x,y))\bigg{)},$
$k_{n}^{\operatorname{sep},q}(x,y)$ $=\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j})\bigg{)}.$

### v4

@@ Abstract
$C^{*}$-algebra-valued kernels could pave the way for the next generation of kernel machines. To further our fundamental understanding of learning with $C^{*}$-algebraic kernels, we propose a new class of positive definite kernels based on the spectral truncation. We focus on kernels whose inputs and outputs are vectors or functions and generalize typical kernels by introducing the noncommutativity of the products appearing in the kernels. The noncommutativity induces interactions along the data function domain. We show that the proposed kernels fill the gap between existing separable and commutative kernels. We also propose a deep learning perspective to obtain a more flexible framework. The flexibility of the proposed class of kernels allows us to go beyond previous separable and commutative kernels, addressing two of the foremost issues regarding learning in vector-valued RKHSs, namely the choice of the kernel and the computational cost.
Keywords: Kernel methods, $C^{*}$-algebra, Positive definite kernel, Ridge regression
@@ 1 Introduction
Kernel methods have been one of the most fundamental tools in machine learning (Schölkopf and Smola, 2001; Gretton et al., 2007; Hofmann et al., 2008; Muandet et al., 2017). They have been applied, for example, to ridge regression, principal component analysis, and support vector machine. Kernel methods are characterized by reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels. Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel. Product kernels, which are constructed by the product of multiple kernels, have also been considered (Schölkopf and Smola, 2001; Thomas, 2008).
@@
Standard positive definite kernels are scalar-valued, and are well-suited to learn scalar-valued functions. Kernel methods for vector- and function-valued outputs have also been investigated (Álvarez et al., 2012; Kadri et al., 2016). The kernels, in these cases, are instead operator-valued, and the associated feature space is vector-valued RKHSs (vvRKHSs) (Kadri et al., 2012, 2016; Minh et al., 2016). There are at least two challenges for vvRKHS methods: the computational cost and choice of kernels. A typical kernel is the separable kernel, which is defined by the product of a scalar-valued kernel and a positive semi-definite operator (Álvarez et al., 2012). Another typical kernel is the commutative kernel, which is defined only with the pointwise calculation of functions or vectors (Hashimoto et al., 2021). Although applying the separable and commutative kernels is computationally efficient, there is a crucial shortcoming for each kernel. Separable kernels identify dependencies between input and output variables separately, and cannot reflect information of input variables properly to output variables. The output is determined only by the global information of the input. On the other hand, commutative kernels only identify the pointwise (completely local) dependencies. Indeed, they are two extreme cases regarding the dependencies between input and output variables. Several attempts have been made to construct kernels that go beyond separable and commutative kernels. A typical nonseparable kernel is the transformable kernels, which is characterized by a map that can incorporate the information of input variables with the output variables. Huusari and Kadri (2021) proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement. Hashimoto et al. (2023a) proposed to use the product of circulant matrices and general squared matrices to construct kernels. Using this kernel, one can generalize the convolution and capture the effect of interactions of different Fourier components. However, if we need an $m$-dimensional vector-valued outputs with these kernels, then we have to construct an $mN$ by $mN$ Gram matrix, where $N$ is the sample size, and the computational cost is $O(m^{3}N^{3})$ in general. Thus, with vvRKHSs, to go beyond separable and commutative kernels, the computational cost is significant.
@@
In this work, we address the two challenges of the computational cost and the choice of kernels by introducing a new class of kernels based on the framework of reproducing kernel Hilbert $C^{*}$-module (RKHM) (Hashimoto et al., 2021). RKHM is a generalization of RKHS by means of $C^{*}$-algebra. $C^{*}$-algebra is a generalization of the space of complex values and has structures of the norm, product, and involution (Murphy, 1990; Lance, 1995). It unifies operators and functions. In this framework, kernels are generalized to $C^{*}$-algebra-valued kernel functions and allow us to consider function-valued kernels, leading to function-valued Gram matrices. By evaluating the values of the function-valued Gram matrix at $m$ different points, we obtain $m$ scalar-valued Gram matrices. This allows us to obtain an $m$-dimensional vector-valued outputs with the computational cost of $O(mN^{3})$, which alleviates the dependency on $m$ from cubic to linear compared to the case of vvRKHSs with nonseparable kernels such as transformable kernels. We summarize the difference between the proposed and existing kernels in Table 1.
@@
To obtain kernels going beyond the separable and commutative kernels with low computational cost, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra (D’Andrea et al., 2014; van Suijlekom, 2021; Connes and van Suijlekom, 2021). The proposed kernels are parameterized by a natural number $n$ corresponding to the dimension of the truncated space. They can be applied to both vector and functional inputs. For vector inputs, we regard the elements of them as the values of functions. We approximate the input functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices, whose $(i,j)$-entry depends only on $i-j$. Thus, $n$ describes the resolution of the discretization, and we call it the truncation parameter. Indeed, $n$ describes the noncommutativity of the kernel. We show that the proposed kernels converge to the commutative kernels as $n$ goes to infinity. On the other hand, if $n=1$, then the proposed kernels are separable kernel. Thus, we can control local and global dependencies through $n$. By introducing the parameter $n$ and setting $n$ to balance the information of gloal dependencies and that of the local dependencies, we obtain higher performance compared to the separable and commutative kernels. We also propose a deep model with the proposed kernels to make the framework more flexible and capture both local and global dependencies of the output on the input.
In the setting of functional data (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021), we have focused on how to interpolate observed data to obtain functions (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021). On the other hand, the proposed truncation kernels shed light on the good effects of discretization (setting $n$ as a finite number) on the learning process.
@@
  • We propose spectral truncation kernels, a new class of function-valued positive definite kernels that go beyond separable and commutative kernels with low computational cost. (see Definition 6)
  • The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter, which adjusts the global and local interactions along the data function domain by introducing the noncommutativity into the learning process. (see Subsection 3.1)
  • We propose a deep approach to make the framework more flexible (see Section 6).
Table 1: Summary of the existing and the proposed kernels
Type of kernels Computational cost (ridge regression) Extraction of local information Extraction of global information
Separable $O(mN^{3})$$\times$ $\checkmark$
Commutative $O(mN^{3})$checkmark $\times$
Transformable $O(m^{3}N^{3})$checkmark $\checkmark$
Proposed (with RKHM) $O(mn^{2}N^{2}+mN^{3})$checkmark $\checkmark$
@@ 2 Preliminaries
We review some mathematical notions and existing studies.
@@ 2.2 Spectral truncation on the torus
As we mentioned in Example 1, the product in $C(\mathbb{T})$ is commutative. However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a matrix, we can obtain a noncommutative product structure. Let $e_{j}$ be the Fourier function defined as $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$. Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square-integrable functions on $\mathbb{T}$. Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$. We approximate $M_{x}$ by $P_{n}M_{x}P_{n}$, i.e., by truncating the spectrum. Then, the $(j,l)$-entry of the representation matrix $R_{n}(x)\in\mathbb{C}^{n\times n}$ of $P_{n}M_{x}P_{n}$ is written as
@@
where $F_{n}(t)=(1/n)\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution. If $n=1$, then $F_{n}(z)=1$, and as $n$ goes to infinity, $F_{n}$ goes to the delta function. More precisely, the following proposition holds (Brandolini and Travaglini, 1997; van Suijlekom, 2021), which implies that for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$.
@@
A generalization of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated (Travaglini, 1994; Brandolini and Travaglini, 1997). Let $P$ be a polyhedron and $jP=\{jz\,\mid\,z\in P\}$ for $j\in\mathbb{Z}$. For $t\in\mathbb{T}^{q}$, the Fejér kernel on $\mathbb{T}^{q}$ is defined as
@@
2.3 Existing operator-valued kernels
We review existing operator-valued kernels and discuss their advantages and shortcomings. We especially focus on the case $\mathcal{X}=\mathcal{A}^{d}$ and the application of kernel ridge regression in this paper. For more details about the operator-valued kernels, see, for example, Álvarez et al. (2012).
Typical existing kernels are summarized as follows: Let $\tilde{k}:\tilde{\mathcal{X}}\times\tilde{\mathcal{X}}\to\mathbb{C}$ be a complex-valued positive definite kernel, $\mu$ be a finite positive measure on $\mathbb{T}^{2}$, and $S:\mathbb{T}\times\mathcal{X}\to\tilde{\mathcal{X}}$. Consider a kernel $k:\mathcal{X}\times\mathcal{X}\to\mathcal{B}(L^{2}(\mathbb{T}))$ defined by
$\left\langle u,k(x,y)v\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}^{2}}u(z)\tilde{k}(S(z,x),S(t,y))v(t)\mathrm{d}\mu(t,z)$ (2)
for $u,v\in L^{2}(\mathbb{T})$.
In general, we have to construct a Gram matrix whose elements are operators. Typically, if we need an output function evaluated at $m$ points, then we discretize the operator on $O(m)$-dimensional space and obtain an $O(mN)$ by $O(mN)$ Gram matrix. Then, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m^{3})$. Therefore, the computational cost is significant. As a special case of the kernel (2), the following two kernels are efficient in the sense of the computational cost.
Separable kernel
If we set $\tilde{\mathcal{X}}=\mathcal{X}$ and $S(t,x)=x$, then the kernel $k$ defined in Eq. (2) is called a separable kernel. In other words, let $A\in\mathcal{B}(L^{2}(\mathbb{T}))$ be a Hermitian positive semi-definite operator defined as $\left\langle u,Av\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}}\overline{u(t)}v(z)\mathrm{d}\mu(z,t)$. The kernel $k:\mathcal{X}\times\mathcal{X}\to\mathcal{B}(L^{2}(\mathbb{T}))$ defined as
$k(x,y)v(z)=\tilde{k}(x,y)Av(z)$
is called a separable kernel.
For separable kernels, the computation is reduced to the computation of the Gram matrix of $\tilde{k}$. Thus, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m)$, where $N$ is the sample size. However, as we can see from the definition, it identifies dependencies between input and output variables separately and cannot reflect local information of input variables to output variables. Indeed, with separable kernels, we cannot specify the relationship between the value of the output function at $z$ and the value $x(z)$ of the input $x$ at $z$ although they have strong connection in many cases. For example, if we have an input $[x_{1},\ldots,x_{d}]\in\mathcal{A}^{d}$ as $d$ time-series explanatory variables (i.e., $x_{1},\ldots,x_{d}$ are functions with respect to time) and try to obtain an output function as a response variable, the values of $x_{1}(z),\ldots,x_{d}(z)$ at time $z$ is strongly related to the value of the output at time $z$. In this case, the separable kernels are not suitable for extracting the relationship between $x_{1}(z),\ldots,x_{d}(z)$ and the values of the output at $z$.
Commutative kernel
[…]
$k(x,y)v(z)=\tilde{k}(x(z),y(z))v(z).$
We call this kernel the commutative kernel.
For commutative kernels, the computation is reduced to the computation at each $z$. Thus, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m)$. However, as we can see from the definition, it only identifies completely local relationship between the input function and the output function. The value of the output function at $z$ is determined only with the value of the input function at $z$. In the same example as separable kernels, the values of $x_{1}(z),\ldots,x_{d}(z)$ at time $z$ is strongly related to the value of the output at time $z$, but may also be related to $y(z+t)$ for $t\in[-T,T]$ for a small number $T$. In this case, the commutative kernels are not suitable for extracting the relationship between $x_{1}(z),\ldots,x_{d}(z)$ and the values of the output around $z$, not only exactly at $z$.
As we have seen, although the separable and commutative kernels are computationally efficient, they are two extreme cases regarding the description of the relationship between the input and output functions; the separable kernels identify dependencies between input and output variables separately (only with global information) and the commutative kernels only identify the commutative (completely local) dependencies. The following existing kernels fill a gap between the above two kernels, but computationally expensive.
Transformable kernel
If we set $\mu$ as the Lebesgue measure, then the kernel $k$ defined in Eq. (2) is called the transformable kernel. In this case, the kernels can identify the dependency between input and output variables through the map $S$. However, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m^{3})$ as we discussed at the beginning of this section. In addition, we have to determine the map $S$, but it is not easy to interpret in general.
Combination of separable and transformable kernels
Considering the sum of separable and transformable kernels and the product of these kernels have also been proposed. However, we have the same shortcomings of separable and transformable kernels.
@@ 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. To obtain vector- or functional-valued outputs, applying vvRKHSs has been investigated. However, as we stated in Section 1, to go beyond the separable kernel and reach higher performance, the computational cost scales as $O(m^{3}N^{3})$ for $m$-dimensional outputs with $N$ samples. (See Appendix 2.3.) To go beyond the separable and commutative kernels with lower computational cost, we consider function-valued ($\mathcal{A}$-valued) kernels and RKHMs. As we will discuss in the last part of Section 5, the computational cost depends on $mN^{3}$ if we use an $\mathcal{A}$-valued kernel.
We first combine existing kernels (Hashimoto et al., 2021, 2023a; Schölkopf and Smola, 2001; Kadri et al., 2012; Álvarez et al., 2012) and obtain the following typical commutative $\mathcal{A}$-valued kernels. Here, we consider the case where the set $\mathcal{X}$ for data is $\mathcal{X}\subseteq\mathcal{A}^{d}$.
@@
    product kernel $k^{\operatorname{prod},q}(x,y)(z)=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z))$.
Here, $q\in\mathbb{N}$ is the degree of the products in the kernel, $\alpha_{i}\geq 0$ is the parameter of $k^{\operatorname{poly},q}$, and for $i=1,2$ and $j=1,\ldots,q$, $\tilde{k}_{i,j}:\mathbb{C}\times\mathbb{C}\to\mathbb{C}$ is a complex-valued continuous positive definite kernel.
We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the typical kernels in Example 2. The kernels in Example 2 involve the product of functions in $\mathcal{A}$. For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1). Thus, as we also discussed in Subsection 2.3, the commutative kernels defined in Example 2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$. If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship. On the other hand, if we introduce the truncation parameter $n$ and transform $x$ and $y$ into Toeplitz matrices using the map $R_{n}$ defined in Subsection 2.2, then the product of $R_{n}(x)$ and $R_{n}(y)$ is noncommutative. Focusing on this feature, we define $\mathcal{A}$-valued kernels based on Example 2, but they have interactions along $z\in\mathbb{T}$.
@@
$k_{n}^{\operatorname{prod},q}(x,y)$ $=S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(\tilde{k}_{1,j}(x,y))^{*}\prod_{j=1}^{q}R_{n}(\tilde{k}_{2,j}(x,y))\bigg{)}.$